OpenStudy (anonymous):

evaluate limit: lim x approaches 0 e^(5x)-1-5x/x^2....

OpenStudy (anonymous):

no i got it to be: 5(e^5x)-1-5x/2x

OpenStudy (anonymous):

Consider only x/x^2 Because the bottom exponent greater, it goes to 0

OpenStudy (anonymous):

i got the answer to be 5/2....

OpenStudy (anonymous):

e^5x=1 So 1-1=0. The lim is 0

OpenStudy (anonymous):

yes...and..

OpenStudy (anonymous):

and..what?

OpenStudy (anonymous):

i said the limit is 0...but i wanna know what the answer is so that i can follow it up...i dont know wher 5/2 come from

OpenStudy (anonymous):

Oh, sorry I was reading the rules for infinity. 0/0 Lhopital

OpenStudy (anonymous):

Lhopital produces 5/x That lim goes to infinity.

OpenStudy (anonymous):

yes i also derivtive x^2 to be 2x...

OpenStudy (anonymous):

5/2x lim is infinity

OpenStudy (anonymous):

ok thanks : )

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ e^{5x} -1 -5x}{x^2} \neq \infty\]

OpenStudy (anonymous):

differentiate top and bottom once

OpenStudy (anonymous):

\[\frac{5e^{5x} -5}{2x} \]

OpenStudy (anonymous):

still 0/0 differentiate again

OpenStudy (anonymous):

\[\frac{25e^{5x}}{2}\]

OpenStudy (anonymous):

now sub x=0

OpenStudy (anonymous):

\[\frac{25}{2}\] is the answer

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I thought only the 5x was over x^2. Use brackets rmalik

OpenStudy (anonymous):

Firstlt, after pluging 0, we get 0/0, there fore, we need to take the derivative of both the denomenator and nominator, then, we will find that it still 0, take the second derivative and you will get 25/2 e^5x-1-5x/x^2 ==>5e^5x-5/2x ==>25e^5x/2 now, plug 0 into the equation and you will get 25/2 !